S214 Experiment for eta/omega experiment

ch2 Target ladder TORCH1 images
dsc00012.jpg dsc00013.jpg dsc00014.jpg dsc00015.jpg dsc00015.jpg dsc00015.jpg dsc00018.jpg dsc00010.jpg dsc00020.jpg dsc00021.jpg dsc00022.jpg

tmc manual proposal for s214 torch test results torch test results part 2 DAQ memo
r118 memo

Q & A for S214 participants
kinematics calculation for deuteron beam energy

Present status

31 Jan 04

run118 memo uploaded, summarizing the achievements and problems.

16 Jan 04

Torch analysis note part 2 uploaded.

13 Jan 04

Torch, DC installed at S2. cables connected.
Efficiency checked for DC. Remote HV control ready. LED pulser being prepared.
Outa arrived.

12 Jan 04

Miki finished installing HV boosters to Torch1 and checked its dark current signals. looks ok. With 4 * 32 ns full range, 700 lams/s achieved for 4 TMC modules. (We need 6 for DC). Start to think about setting up two independent DAQ system.
Naoya recovered from cold.
Torch beam test analysis report is now available. The position resolution of 1.2 mm (sigma) has been achieved for 3He (Ryu&Miki).

11 Jan 04

I found that reading TMC modules slows down the DAQ. Even down to 100 lam/s with half of necessary number of channels...

10 Jan 04

DC moved to a table at S2.
Torch1 check being done. booster cable connection under construction.
DAQ being setup. 3500 lam/s w. 32 words achieved. 2500 lam/s w. about 64 word.
TMC being tested. looks ok.

9 Jan 04

DAQ setup being done.
it still takes time to read modules and we still cannot make a test of the TMC.
Gas system for DC being prepared by Ken Kalle and Irina.
PMTs installed for TORCH1.

8 Jan 04

We found that the target #49 PE 1g is damaged. A brown spot was observed at the center. We started to replace the target however the SIS (incl FRS TA) area will be closed on the 12/Jan(Mon) and it is unclear if the target will be ready before.
Booster cables are taken care for Torch1 by Miki and Naoya.
Borrowed rio3 based camac controller from Nik.
Kalle and Irina arrived.

7 Jan 04

Ken arrived.

6 Jan 04

Boxes from KEK arrived.

5 Jan 04

Miki Naoya Kenta arrived. Cloudy sky with snow. very cold to the bone.

Our test beamtime for optics/detector calibration will be between 29th/Nov and 3rd/Dec. We need 9 modules of 16ch ECL differential input (34pin flat cable) TDC modules.

Status Log

6 Nov

* DC
We setup the gas system for drift chamber(XX', XX') conditioning and flow P10 gas into them.
* We took all necessities that we put backside of S2 out of the area.

7 Nov

We supply HV to DC for aging. Now HV=900V/850V

* We check the modules inside of the shelf of Messhutte. 4 modules of PHILLIPS 7186H are inside. And I check the shelf on the roof at S2. 1 module of PHILLIPS 7186H we have.
So now we have 5 modules of 7186H. We need more than 7 modules of that.

We start re-making of TORCH(eg. making PMT-booster-system).

8 Nov

aging. Now HV=1150/1100V
Naoya and Kenji finished making connectors for booster system.
We remove all PMT's from TORCH.

9 Nov

* Ken and Kenta arrived at GSI.
We are making booster system of TORCH.
* DC
aging. Now HV=1100/1150V

10 Nov

Although Ken, Naoya and Kneta tried to test the DC efficiency, It's very noisy.
Now HV=1050/1100V
Kenji and I were making booster system.

11 Nov

Now HV=1050/1100V
Noise is reduced with the HV filter (Threshold is set about 80mV)
Ken, Naoya and Kenta will start checking the efficiency with a scintillator tomorrow.

The PMT's which were exported from Japan were arrived !
Then we are ready for re-making TORCH.
Kenji and I were making booster system.

Area will be closed tomorrow or the day after tomorrow.

*Present status at S4

12 Nov

The modules that are exported from Sweden are arrived. The modules should be returned in the end of this test experiment (middle of December).
Tomorrw or the day after tomorrow, we can start testing the efficiency with Ar and Isobutaine gas (We need to make the gas system).
We finished making the booster system except for a feedthrough system.
* Support Structure at S4
We'll try to talk to Adolf about that tomorrw, though he is now very busy.
Kenta start setup DAQ.
Ken, Naoya and Ken assemble the CRATE (CAMAC and VME) in front of the gate at S4.
* Area
Area will be closed on Friday.
*Present status at S4

13 Nov

* DC
Ken and Naoya made gas system(Ar + Isobutaine). They'll setup the Mass Flow Controller.
Kenji and I tried to sort the PMT gain with LED pulser. We'll install or put the new PMT in TORCH-box once, and sort the PMTs by the single photon peak spectra with driving LED puser.
DAQ and NIM crates are assembled in front of the gate at S4. Kenta are making DAQ system.
* First Spectra
Kenji and I setup the scintillator for testing DAQ system. And we can see the QDC spectra(90Sr) online. This is the first spectra....
* Memo
All our packs that should be arrived at GSI are recieved. Thanks !

14 Nov

* DC
Naoya, Kenji and Kenta exchange gas for DC from P10 to Isobutane + Ar.
They set up Mass Flow Controller (3:1).
Now HV = 0V
Tomorrow we'll check the efficiency.
Kenji and I put 8 new-PMTs into the TORCH-box. We drive the LED puler and take QDC data. We can see single photon peaks.

15-19 Nov

Naoya, Kenji and Ken prepare the BNC long cables for TORCH and Scintillators. Because we setup the DAQ system in front of the gate at S4, we need about 20m length BNC cabls from S4 to CAMAC CRATE. All cables are ready except for the flat cabels for DC. We'll prepare the HV cables in a few days.
We check the efficiency for all DC planes(1-XX', 2-XX').
In case that the threshold is set at 100mV, the efficiency is
1-X= 98.5% (HV = 1100/1050V)
1-X'= 97.8% (HV = 1100/1050V)
2-X= 98.0% (HV =1100/1050V)
2-X'= 96.6%(HV = 1100/1050V).
Today we mount the DC(1-XX',2-XX') on the support structure.
Now HV=off.
We ask Karl Heints to make hole at the back steel of TORCH box. We'll make the booster connectors feedthrough system with the hole.
I took the QDC data by LED pulser. I cannot understand the pedestal thickness. Very very thin. I'll check again tomorrow.
Kenta and I talked to Adolf how to do for the S4 supprot structure.

20 Nov

We made a panel for booster feedthrough of TORCH. We'll check the light leak tomorrow.
*DC, Gas
We tried to controll "Mass Flow Controller" properly, because the displayed value of the blending ratio is wrong. Now it's the proper value.

21 Nov

We sort the new PMT's gain with the HV is 1600V. I cannot understand why the pedestals seem to be thin (1-2 bin). I'll check again.
photo 1
* DC
Naoya, Kenji and Ken mount the DC on the support structure. Now HV=950/1000V. photo 2
photo 3
photo 4
photo 5
photo 6
photo 7
photo 8
photo 9

22 Nov

Ken and I checked the QDC spectra (HV=1600V). We cannot understand
Kenji and Naoya put the connectors(34pin) to the Flat cables for DC.

23-24 Nov

Naoya and Kenji checked dead channels of flat cables and the length.
Ken tried to read the data by PHILLIPS 7186H TDC with Time Calibrator.
I measured the linearity of QDC modules what we have. There exist the linear corelation of channels of the same modules, but there is no corelation between the different modules. Then we have to use the TECHNO-16CHANNEL QDC for TORCH.
Kenta and I calculated the number of channels for the single photon peak. Consequentry the pulse height of the single photon should be about 25 mV.
Then we can take the 3He data without OVER FLOW and without PMT Amp.(TECHNO QDC:12bit 4096CHANNEL)

Eberhard arrived at GSI today !

Agnieszka will arrive 26th of Nov (Wed).

Ryu will arrive 27th of Nov.

25 Nov

Kenji and Ken prepared the flat cables.
They read the data of the amp-distriminator test input with PHILLIPS 7186H TDC.
Naoya, Kenta and I installed new PMTs in TORCH. The order of the PMT is random to the PMT gain. Now we take the LED Data without booster.

* Support Structure
Kenta and I talked about it to Adolf. Readout list 4300B manual

29 Nov

To Shift crew A of 30/Nov.
Summary of shift B of 29/Nov.:
We found that the MW21 X cables were connected incorrectly. This caused wrong direction of horizontal axis in the MW21 spectrum. Therefore we started again from the offset measurement of MW21 by the S2 slits.
Also, we found that the S1 position is shifted by 4 mm. This was measured by the MW11 and by the S1 slits. Thus we scaled TA-S1 and S1-S2 separately and now S1 position looks ok.
However, at S2, MW22 spectrum is well centered, MW21 shows -8 mm. We couldn't understand the reason even now.
We proceeded to S3 and S4 and they are centered.
Then, we started to measure the effective thickness of the S2 degrader as written in the log book.
In the morning of 30th, I think it is a good idea to start with dispersion and focal plane measurement. Scaling FRS and changing incident angles is necessary for this.